# UU-Mamba: Uncertainty-aware U-Mamba for Cardiovascular Segmentation

### v1

@@ 4.3 Ablation Study
We perform an ablation study to investigate the impact of integrating Sharpness-Aware Minimization (SAM) optimization [13] into our UU-Mamba model, alongside traditional and uncertainty-aware loss functions. This study spans three datasets: ACDC [2], Aorta [10, 11], and ImageCAS [9], and employs 3D loss surface visualizations [27] using ParaView [64, 65, 66] to demonstrate the smoother loss landscapes indicative of model robustness and improved generalization.
4.3.1 Impact of Traditional and Uncertainty-Aware Loss Functions
@@
In the ACDC dataset [2], the model trained with only CE loss achieves a Dice Similarity Coefficient (DSC) of 92.263%. When incorporating the uncertainty-aware loss, the DSC improves to 92.602%. This enhancement reflects the efficacy of the uncertainty-aware approach in improving segmentation resilience by carefully managing prediction uncertainty and refining the model’s focus on areas of high confidence, as also shown in Table 2.
@@
For the Aorta dataset [10, 11], training with only CE loss yields an average Dice Similarity Coefficient (DSC) of 73.761%. Incorporating the uncertainty-aware loss results in notable improvements across various anatomical zones and arteries, raising the average DSC to 75.053% and average NSD to 91.747%, also reducing the Mean Squared Error (MSE). These improvements, as shown in Table 3, highlight the uncertainty-aware loss’s ability to handle complex segmentation tasks by balancing confident predictions with uncertain areas.
@@
For the ImageCAS dataset [9], the model trained with only CE loss achieves an average Dice Similarity Coefficient (DSC) of 79.496%. By incorporating the uncertainty-aware loss, the average DSC improves to 81.146% and NSD improves to 88.045%, demonstrating the advantage of this comprehensive loss function in enhancing segmentation accuracy, particularly in challenging regions, as shown in Table 4.
4.3.2 Enhancements with Sharpness-Aware Minimization optimization
@@
Building on the incorporation of uncertainty-aware loss, we further integrate Sharpness-Aware Minimization (SAM) optimization [13] to explore its additional benefits. SAM is designed to steer the training process toward flatter minima, which are associated with improved generalization in neural network models. Figure 6 illustrates a comparison of the loss landscapes between models trained with and without SAM optimization.
@@
In Table 2, incorporating SAM with the uncertainty-aware loss increases the ACDC dataset’s DSC to 92.787%, the highest among the tested methods, thus validating SAM’s role in enhancing segmentation precision and generalization. In the Figure 6, the loss landscape on the ACDC dataset is shown for models with and without SAM optimization. The 3D loss surface plot of the model without SAM optimization exhibits a broader range of loss values, characterized by sharper and more erratic loss contours. In contrast, the model utilizing SAM optimization displays a flatter and smoother loss landscape, indicating improved stability and generalization.
@@
As shown in Table 3, incorporating SAM into the model training process increases the average DSC to 77.084%, along with significant improvements in both DSC and NSD metrics. SAM optimization leads to smoother and more stable loss surfaces, as demonstrated in Figure 6. Without SAM, the 3D loss surface exhibits sharper and more rugged contours, particularly along the edges. In contrast, the model with SAM displays a flatter, more stable loss landscape, indicating improved robustness and generalization.
@@
Notably, the UU-Mamba model achieves the highest DSC scores in 17 out of 24 anatomical regions and the highest NSD values in 13 out of 24 regions, as detailed in Table 5. These results underscore the superior generalizability and accuracy of the UU-Mamba model. The consistent top performance across most regions highlights the significant benefits of combining uncertainty-aware loss with SAM optimization to enhance segmentation outcomes.
@@
As detailed in Table 4, SAM optimization pushes the ImageCAS dataset performance metrics to the highest levels observed in this study. In Figure 6, the 3D loss surface of the model without SAM shows greater variability, especially in the bottom right region. By contrast, with SAM optimization, the loss surface becomes much smoother, indicating improved model consistency, robustness, and generalization across diverse cardiovascular imaging scenarios.
@@
4.4 Robustness Analysis
We perform experiments to evaluate each method on the Mean Squared Error (MSE) of the DSC scores to assess their robustness quantitatively. The MSE is calculated as shown in Eq. (10). Results are shown in the Tables 2, 3, and 4. These results show that the uncertainty-aware loss reduces the MSE compared to the standard CE loss, reflecting its ability to better address the variability and uncertainty in the data. The inclusion of SAM optimization significantly decreases the MSE, achieving the lowest error value. This reduction in MSE highlights SAM’s role in minimizing errors and producing more accurate segmentation maps.
@@ 4.4 Robustness Analysis
5 Conclusion
We present a novel model, UU-Mamba, specifically developed for the purpose of segmenting cardiovascular MRI and CT data. This model combines the U-Mamba architecture with an uncertainty-aware loss function and the SAM optimizer, resulting in a substantial enhancement of biological picture segmentation. It achieves improved generalization and boundary accuracy. The uncertainty-aware loss function integrates region-based, distribution-based, and pixel-based losses to enhance segmentation performance by effectively managing jobs and prioritizing confident predictions. Simultaneously, the SAM optimizer directs the model towards flat minima in the loss landscape, improving its ability to withstand challenges and decreasing the likelihood of overfitting, ultimately resulting in more accurate segmentation. Aside from doing our main tests on the ACDC dataset [2], we also assessed the performance of UU-Mamba on two other datasets: ImageCAS [9] and Aorta [10, 11]. The model scored the greatest average DSC, NSD, and MSE on the ImageCAS dataset, demonstrating a considerable improvement compared to the baseline models. UU-Mamba demonstrated superior performance compared to other models on the Aorta dataset, earning the greatest average DSC in 17 out of 24 anatomical regions and the highest average NSD in 13 out of 24 anatomical regions. The data illustrate that the model is highly effective in different anatomical regions and segmentation tasks, highlighting its versatility and strength in numerous medical imaging situations. The comparative analysis conducted on five prominent models establishes the superiority of UU-Mamba. It achieves a DSC of 92.787% on the ACDC dataset, demonstrating high accuracy and robustness in segmenting various datasets, such as ImageCAS and Aorta.
@@ 5 Conclusion
Acknowledgments. This work is supported by the University at Albany Start-up Grant. The authors appreciate the computational resource provided by the University at Albany – SUNY.

### v2

@@ IV-D Ablation Study
We perform an ablation study to investigate the impact of integrating Sharpness-Aware Minimization (SAM) optimization [foret2021sharpness] into our UU-Mamba model, alongside traditional and uncertainty-aware loss functions. This study spans three datasets: ACDC [bernard2018deep], Aorta [imran2024cis, krebs2024volumetric], and ImageCAS [zeng2023imagecas], and employs 3D loss surface visualizations [Hao2018] using ParaView [ahrens2005ParaView, ayachit2015, ayachit2021] to demonstrate the smoother loss landscapes indicative of model robustness and improved generalization.
IV-D1 Impact of Traditional and Uncertainty-Aware Loss Functions
@@
In the ACDC dataset [bernard2018deep], the model trained with only CE loss achieves a Dice Similarity Coefficient (DSC) of 92.263%. When incorporating the uncertainty-aware loss, the DSC improves to 92.602%. This enhancement reflects the efficacy of the uncertainty-aware approach in improving segmentation resilience by carefully managing prediction uncertainty and refining the model’s focus on areas of high confidence, as also shown in Table II.
@@
For the Aorta dataset [imran2024cis, krebs2024volumetric], training with only CE loss yields an average Dice Similarity Coefficient (DSC) of 73.761%. Incorporating the uncertainty-aware loss results in notable improvements across various anatomical zones and arteries, raising the average DSC to 75.053% and average NSD to 91.747%, also reducing the Mean Squared Error (MSE). These improvements, as shown in Table III, highlight the uncertainty-aware loss’s ability to handle complex segmentation tasks by balancing confident predictions with uncertain areas.
@@
For the ImageCAS dataset [zeng2023imagecas], the model trained with only CE loss achieves an average Dice Similarity Coefficient (DSC) of 79.496%. By incorporating the uncertainty-aware loss, the average DSC improves to 81.146% and NSD improves to 88.045%, demonstrating the advantage of this comprehensive loss function in enhancing segmentation accuracy, particularly in challenging regions, as shown in Table IV.
IV-D2 Enhancements with Sharpness-Aware Minimization optimization
@@
Building on the incorporation of uncertainty-aware loss, we further integrate Sharpness-Aware Minimization (SAM) optimization [foret2021sharpness] to explore its additional benefits. SAM is designed to steer the training process toward flatter minima, which are associated with improved generalization in neural network models. Figure 6 illustrates a comparison of the loss landscapes between models trained with and without SAM optimization.
@@
In Table II, incorporating SAM with the uncertainty-aware loss increases the ACDC dataset’s DSC to 92.787%, the highest among the tested methods, thus validating SAM’s role in enhancing segmentation precision and generalization. In the Figure 6, the loss landscape on the ACDC dataset is shown for models with and without SAM optimization. The 3D loss surface plot of the model without SAM optimization exhibits a broader range of loss values, characterized by sharper and more erratic loss contours. In contrast, the model utilizing SAM optimization displays a flatter and smoother loss landscape, indicating improved stability and generalization.
@@
As shown in Table III, incorporating SAM into the model training process increases the average DSC to 77.084%, along with significant improvements in both DSC and NSD metrics. SAM optimization leads to smoother and more stable loss surfaces, as demonstrated in Figure 6. Without SAM, the 3D loss surface exhibits sharper and more rugged contours, particularly along the edges. In contrast, the model with SAM displays a flatter, more stable loss landscape, indicating improved robustness and generalization.
@@
Notably, the UU-Mamba model achieves the highest DSC scores in 17 out of 24 anatomical regions and the highest NSD values in 13 out of 24 regions, as detailed in Table V. These results underscore the superior generalizability and accuracy of the UU-Mamba model. The consistent top performance across most regions highlights the significant benefits of combining uncertainty-aware loss with SAM optimization to enhance segmentation outcomes.
@@
As detailed in Table IV, SAM optimization pushes the ImageCAS dataset performance metrics to the highest levels observed in this study. In Figure 6, the 3D loss surface of the model without SAM shows greater variability, especially in the bottom right region. By contrast, with SAM optimization, the loss surface becomes much smoother, indicating improved model consistency, robustness, and generalization across diverse cardiovascular imaging scenarios.
@@
IV-E Robustness Analysis
We perform experiments to evaluate each method on the Mean Squared Error (MSE) of the DSC scores to assess their robustness quantitatively. The MSE is calculated as shown in Eq. (10). Results are shown in the Tables II, III, and IV. These results show that the uncertainty-aware loss reduces the MSE compared to the standard CE loss, reflecting its ability to better address the variability and uncertainty in the data. The inclusion of SAM optimization significantly decreases the MSE, achieving the lowest error value. This reduction in MSE highlights SAM’s role in minimizing errors and producing more accurate segmentation maps.
@@ IV-E Robustness Analysis
V Conclusion
We present a novel model, UU-Mamba, specifically developed for the purpose of segmenting cardiovascular MRI and CT data. This model combines the U-Mamba architecture with an uncertainty-aware loss function and the SAM optimizer, resulting in a substantial enhancement of biological picture segmentation. It achieves improved generalization and boundary accuracy. The uncertainty-aware loss function integrates region-based, distribution-based, and pixel-based losses to enhance segmentation performance by effectively managing jobs and prioritizing confident predictions. Simultaneously, the SAM optimizer directs the model towards flat minima in the loss landscape, improving its ability to withstand challenges and decreasing the likelihood of overfitting, ultimately resulting in more accurate segmentation.
Aside from doing our main tests on the ACDC dataset [bernard2018deep], we also assessed the performance of UU-Mamba on two other datasets: ImageCAS [zeng2023imagecas] and Aorta [imran2024cis, krebs2024volumetric]. The model scored the greatest average DSC, NSD, and MSE on the ImageCAS dataset, demonstrating a considerable improvement compared to the baseline models. UU-Mamba demonstrated superior performance compared to other models on the Aorta dataset, earning the greatest average DSC in 17 out of 24 anatomical regions and the highest average NSD in 13 out of 24 anatomical regions. The data illustrate that the model is highly effective in different anatomical regions and segmentation tasks, highlighting its versatility and strength in numerous medical imaging situations. The comparative analysis conducted on five prominent models establishes the superiority of UU-Mamba. It achieves a DSC of 92.787% on the ACDC dataset, demonstrating high accuracy and robustness in segmenting various datasets, such as ImageCAS and Aorta.